Finding the equations of a line using slope-intercept method
Please write the equation of the ff. with the given slope and y-intercept:
1)m=-2 b=0
2)m=2/5 b=-4
3)m=3/5 b=2/3
Pls. reply as soon as possible because i'm still confused...

Answer by CharlesG2(834)

Finding the equations of a line using slope-intercept method
Please write the equation of the ff. with the given slope and y-intercept:
1)m=-2 b=0
2)m=2/5 b=-4
3)m=3/5 b=2/3
Pls. reply as soon as possible because i'm still confused...

not sure what "ff." means maybe it means "the following", but I can show you the equations of those 3 lines
slope-intercept form of a line is y=mx+b where m is the slope or the rise/run, and b is the y-intercept (vertical-intercept or point (0,b))
1. m = -2, b = 0 --> y = -2x
2. m = 2/5, b = -4 --> y = (2/5)x - 4
3. m = 3/5, b = 2/3 --> y = (3/5)x + 2/3